OpenStudy (anonymous):

Salina currently has an account balance of $1,047.69. Her initial deposit on the account was $630 and it earned 3.9% simple interest. How long has Salina held the account?

OpenStudy (jdoe0001):

her current balance is $1,047.69 her initial deposit was $630 the difference is 1047.69 - 630 = 417.69 she earns 3.9% simple interest rate, which the earned interest is 417.69 so 3.9% for this many years has earned $417.69 \(\bf interest = principal(\textit{deposit})\times rate(\textit{in decimal})\times time(\textit{years})\\ \quad \\ i = prt\implies \cfrac{i}{pr}=t\\ \quad \\ 3.9\%\textit{ in decimals is }0.039\\ \quad \\ \cfrac{i}{pr}=t\implies \cfrac{417.69}{630\times 0.039}=t\)

OpenStudy (anonymous):

0.025857

OpenStudy (jdoe0001):

hmmm ? 0.025857?

OpenStudy (jdoe0001):

recheck... again.... seems your PEMDAS may be a bit off

OpenStudy (anonymous):

17

OpenStudy (jdoe0001):

:)

OpenStudy (jdoe0001):

17years

OpenStudy (anonymous):

thanks

OpenStudy (jdoe0001):

yw